OpenStudy (anonymous):

For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. x 8, 16, 24 - y 11, 22, 33 choose an answer. a. yes; y = 2.75x b. yes; y = 0.6875x c. yes; y = 1.375x d. no; y does not vary directly with x

OpenStudy (anonymous):

To see if they vary directly, you can divide each of the y values with the corresponding x value.

OpenStudy (anonymous):

They all are the same number when I divide them.

OpenStudy (btaylor):

yes, since as x increases y does as well. the rate of change is (22-11)/(16-8) = 11/8 = 1.375 This means the slope is 1.375, so C is the answer.

OpenStudy (anonymous):

That number is the direct variation.

OpenStudy (anonymous):

Happy to help:)

OpenStudy (btaylor):

yw,

OpenStudy (anonymous):

If you can help me with this one also, It would be greatly appreciated. :)

OpenStudy (btaylor):

sure

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = -6. What direct variation equation relates x and y? What is the value of y when x = -2? Choose one answer. a. y = -0.75x; 1.50 b. y = -1.33x; 2.67 c. y = 1.33x; -2.67 d. y = 0.13x; -0.25

OpenStudy (btaylor):

if there is a direct relationship, then as x increases, so does y. This implies a positive slope, so we can eliminate A and B immediately.

OpenStudy (btaylor):

Now, we know that the y-value has to increase from when x=-6 to when x = -2.

OpenStudy (btaylor):

It doesn't look like any of the y-values are greater than 8, so I'm kinda stuck...sorry.